OpenStudy (anonymous):

Consider the function z = f(x,y) = 1+sqrt(9+x-y^2) . Find the range of f(x,y). Hint: what are the possible outcomes (z values).

OpenStudy (anonymous):

\[z = f(x,y) = 1+\sqrt(9+x-y^2) \] this?

OpenStudy (anonymous):

it's \[\sqrt{9+x-y^2}\]